# Vibravox: A Dataset of French Speech Captured with Body-conduction Audio Sensors

### v3

@@ III-B Textual data
The text utilized for participant readings originates from the French Wikipedia subset of Common Voice [60]. This textual data was further filtered to produce a simplified dataset with a minimum number of textual and phonemic symbols and to minimize pronunciation uncertainty. To create this simplified dataset, we applied textual filters that excluded most proper names, retaining only common first names, French town or region names, and country names. Moreover, utterances containing numbers, Greek letters, mathematical symbols, or syntactic errors were carefully removed, resulting in a more straightforward dataset that is suitable for accurate pronunciation analysis. Once the unwanted sentences were removed, we applied common text normalization operations, such as lowercasing text and removing ponctuation except for apostrophe. From this normalized text, we also generated the corresponding phonetic transcription using the International Phonetic Alphabet (IPA), which contains 33 phonemes in French. This operation was done with the text phonemizer of the CoML lab [61]. For some words, the phonemizer produced some language switches that we wanted to avoid in order to preserve the minimal French phonetic alphabet. A phase of manual phonetic transcription was then necessary for these unrecognized words, such as ”algorithme”.
@@ III-D Recording protocol
  • speechless-noisy: For a period of 2 minutes and 24 seconds, the subject is required to remain silent, yet free to move, swallow and breathe naturally in a noisy environment created from the AudioSet samples described in subsection III-C. Those samples are played in a spatialization sphere equipped with 56 loudspeakers surrounding the subject. The objective of this phase is to gather realistic background noises that will be combined with the speech-clean recordings to maintain a clean reference.
@@ III-E Post processing
The filtering process comprises two main stages. In the first stage, indicators for evaluating audio capture quality are extracted (Figure 2(a)). In the second stage, these indicators are used to derive a set of filters (Figure 2(b)). The whisper-timestamped model333available at https://github.com/linto-ai/whisper-timestamped constitutes a key component of our indicator extraction process. This model is a speech-to-text engine based on Whisper [29] that applies dynamic time warping [65] to cross-attention weights. By aligning the transcription with the raw waveform, accurate prediction of words timestamps in speech segments can be used to identify sequences containing vocalizations. The indicators we are using are listed below:
  • $\mathrm{PER}[j]$: Phoneme Error Rate are computed for each sample $j\in[1,J]$ using the Levenshtein distance between the phonetic labeled transcription and the phonetic transcription derived from Whisper’s output. Note that we used the exact same phonemization procedure as described in subsection III-C.
@@ III-F Signal processing analysis
Figure 3: Coherence function of body-conducted sensors in quiet conditions
Each sensor reveals distinct bandwidth size that the figure legend follows from left to right in decreasing order. The temple sensor exhibits the smallest bandwidth, barely recording any speech signal above 1500 Hz. The laryngophone follows, with coherence dropping below 0.25 at 2000 Hz. The soft and rigid in-ear sensors exhibit larger bandwidths, yet display clear antiresonance frequencies. The soft in-ear bandwidth is slightly larger than its rigid in-ear counterpart, which may be due to a less effective acoustic seal. Finally, the forehead accelerometer offers the largest bandwidth, yet does not effectively filter out external noise pollution. As discussed in Section IV, this bandwidth is closely correlated with the performance of deep learning models on speech intelligibility, quality, speaker verification, and speech-to-phonemes.
@@ IV Using Vibravox
The Vibravox dataset offers considerable potential across various applications, particularly in improving communication in noisy environments. The speech-noisy subset is of particular interest for showcasing the noise-resilient capabilities of BCMs. However, this subset primarily serves testing purposes as it lacks a clean reference under such conditions. As a consequence, the majority of the dataset is contained within the speech-clean subset, where the airborne microphone serves as the reference. In order to simulate real-world noisy scenarios while retaining a clean reference, data augmentation techniques can be applied by adding audio segments of the speechless-noisy subset with BCM utterances. Furthermore, the Vibravox dataset displays potential for applications such as developing voice commands and monitoring vital signs, including heart rate, particularly when utilising the speechless-clean subset. In the following, we propose three classic tasks in automatic speech processing to investigate the distinctive characteristics of BCMs, establishing baselines for each task.
@@ IV-A Speech Enhancement
The high impedance of BCMs enables them to focus on the wearer’s voice by rejecting ambient noise, but this also affects their ability to capture mid and high-frequency components due to the intrinsic low-pass characteristics of the biological pathway of vibrations between the vocal chords and the body conducted audio sensors. The absence of these frequencies in captured speech directly impacts the perceived quality, intelligibility, and distinctiveness. This highlights the necessity for a post-capture enhancement algorithm.
@@ IV-A1 Model
We adopted the configurable EBEN [17] approach, enriched with recent trends in the audio field. This GAN [68] architecture offers several advantages. In particular, the lightweight design of the 1D U-net-like generator, which is compatible with integration in mobile and wearable systems where memory is limited and real-time is a priority. Additionally, the PQMF bank [69] offers the advantage of manipulating frequency content without the need to consider phase or manage complex values, as it operates directly on waveform signals. It is important to note that the limitations of GANs, including the difficulty of reaching a Nash equilibrium, the lengthy training period, and the additional computational cost associated with the discriminators, do not affect the model once it has been deployed.
@@ IV-A2 Architecture configuration
Separate EBEN models were trained for each sensor in the Vibravox dataset, using sensor-specific hyperparameters detailed in Table IV. The original number of subbands was maintained to a value of $M=4$ for each sensor due to the significant variability in intra-sensor bandwidth among participants, rendering precise spectrogram slicing impractical. Furthermore, a smaller number of subbands allows for a more permissive design of the passband filters with a softer slope for the same resulting aliasing, thus shortening the PQMF kernel length, which is useful for reducing the algorithmic latency. The number of subbands provided to the model, denoted as $P$, was tailored to match the bandwidth of each sensor. For example, the temple vibration pickup captures signals within the 0-2 kHz range, while the forehead accelerometer exhibits good coherence with the reference microphone up to 8 kHz. Consequently, after resampling the signals to 16 kHz for training, only the first band was retained for the temple vibration pickup, while all bands were retained for the accelerometer. In addition, the last configurable parameter, $Q$, which represents the number of bands fed to the PQMF discriminator, has been modified to match $M$ at 4 for all models. It was observed that even when the audio sensors capture information within a specific frequency range, including the corresponding subbands in the discriminator input is beneficial for denoising minor artifacts present in these initial bands. Also, in the case of speech captured in noisy conditions, despite the fact that body conduction sensors are robust to external acoustic pollution, some noise is still captured, providing an additional reason to focus discriminators work on every subband. This minor modification necessitated an adjustment to the number of channels in the PQMF discriminators, in order to align with the divisibility constraints of the grouped convolution.
Figure 4: Spectrograms of signals recorded by the different Vibravox audio sensors in the speech-clean subset and their corresponding EBEN enhanced version
@@ IV-A3 Training procedure
In comparison to the original implementation proposed in [17], we have incorporated a spectral loss $\mathcal{L}_{\mathcal{G}}^{spec}$ into the generator objective. This modification proved to be effective in stabilizing the training process. This loss incorporates the same FFT, hop size, and window length parameters as described in [70]. Implementation is taken from [71]. In contrast to some speech enhancement papers [72, 73], we did not include a reconstructive loss in the temporal domain due to the peculiarities of body-conduction speech. In fact, complex phase shifts occur, and attempting to align the signals to the reference at the sample level degraded performance. We retained the original EBEN implementation’s use of the other two adversarial and a feature matching losses, respectively noted $\mathcal{L}_{\mathcal{G}}^{adv}$ and $\mathcal{L}_{\mathcal{G}}^{feat}$. Also note that a slight adjustment to $\mathcal{L}_{\mathcal{G}}^{feat}$ has been done by normalizing it using the enhanced feature norm. The training strategy remained consistent across speech-clean and speech-noisy experiments, with a constant learning rate of $3e^{-4}$. To ensure reproducibility of experiments, the code has been made available on the GitHub repository https://github.com/jhauret/vibravox, which makes use of the following libraries [74, 75, 76, 77].
@@ IV-A4 Experiments in quiet conditions
For this initial set of experiments, the goal is to evaluate the performance of bandwidth extension in quiet conditions by training EBEN models on the speech-clean subset. A total of 500 epochs were required for convergence on the Vibravox dataset. To address the relatively small amount of data, light data augmentation was applied, including speed perturbation, pitch shifting, and time masking, all implemented in the Torchaudio library [78]. The correlation analysis presented in the configurable EBEN article [17] revealed that STOI [79, 80] and Noresqa-MOS (N-MOS) [81] were the metrics with the highest correlation with MUSHRA studies in the context of speech enhancement on BCMs simulated data. Consequently, these two metrics were retained for inclusion in Table IV which presents the results of the bandwidth extension for each sensor in quiet conditions. The results reveal a clear relationship between the metrics value and the sensors’ available bandwidth. Yet, both STOI and N-MOS performance appear to hit a glass ceiling, plateauing at 0.88 and 4.3 respectively. This saturation could stem from the challenge of extracting fine details from body-conducted signals. While striving to produce authentic samples, the generator only picks one plausible clean signal out of all the possible ones, typical of one-to-many problems [82]. The spectrograms in Figure 4 provide a visual representation of the processing.
@@
Sensor Configuration STOI Noresqua-MOS
@@ IV-A5 Experiments in noisy conditions
After analyzing speech enhancement in quiet conditions, the next focus is a joint denoising and bandwidth extension task for recordings in noisy environments. To this end, the model is tested on the speech-noisy subset; however, this subset cannot be used for training due to the absence of a clean reference. For training, the speechless-noisy subset is leveraged and dynamically mixed with the speech-clean subset. Several configurations were tested, including retraining the EBEN models from scratch for 500 epochs, or initializing from the previously trained speech-clean models, which converged after no more than 200 epochs for microphones with the largest domain gap. Additionally, speech-clean models were directly tested on speech-noisy. Hyperparameters of the training strategy were kept consistent with quiet condition experiments, with the same values for M, P, and Q across all microphones.
Prior to testing on speech-noisy, the model performance was first evaluated on the validation and test splits of the synthetically mixed speech-clean+speechless-noisy data. When testing on speech-noisy, no significant differences were observed in the Noresqua-MOS scores, reflecting a realistic modeling of the training data. When using speech-noisy for evaluating the models, the STOI – which is a reference-based metric – was replaced with the reference-free TorchSquim-STOI [83], which provides an estimate of the same quantity. Although direct comparison between STOI and TorchSquim-STOI is not possible, both metrics showed improvement during training, and the TorchSquim-STOI results remained consistent across both synthetic mixed and real noisy data. The final test results on speech-noisy are presented in Table V.
@@
Sensor Initialization Squim-STOI Noresqua-MOS
@@
The results reveal that microphones with the lowest signal-to-external-noise ratios benefited the most from the retrained strategies. Some microphones, such as the temple vibration pickup, were so insensitive to noise that the non-retrained speech-clean models outperformed them. The Noresqua-MOS scores for the raw signals also increased compared to the speech-clean subset. However, these Noresqua-MOS values cannot be directly compared, as in the quiet conditions, the metric was computed using matched signals, whereas in the noisy conditions, a random headset signal from the speech-clean was used for the calculation.
@@
As with the experiments conducted in quiet conditions, several spectrograms are presented in Figure 5. Synthetically mixed data were selected to visually assess the effect of the models, as both the noise and the clean target signals remain available. This figure highlights several noteworthy observations. Notably, the excessive noise detected in the soft in-ear microphone suggests an insertion defect for this particular sample of speechless-noisy, whereas the rigid microphone appears to provide better acoustic sealing. However, such defects occur with similar frequency in both microphone types and are absent in the majority of cases. Additionally, EBEN models trained for bandwidth extension in quiet conditions exhibit limited noise suppression. In contrast, as shown in line (c), noise-robust EBEN models effectively suppress noise, with the difference particularly pronounced for forehead and in-ear microphones. For throat and temple microphones, which are inherently more resistant to noise, the domain gap between quiet and noisy conditions remains small, enabling the speech-clean EBEN models to maintain strong performance. Achieving comparable denoising performance from the noisy headset using conventional denoising networks would be highly challenging.
@@ IV-B2 Experiments in quiet conditions
For this first set of experiments, we finetuned the wav2vec2.0 models with a constant learning rate of $1e^{-5}$ on the speech-clean subset of the Vibravox dataset for 10 epochs, minimizing the CTC loss [86]. We then tested each fine-tuned model not only on its corresponding test set but also on every other audio sensors, with the PER (Phoneme Error Rate) results shown in Figure 6. As expected, the matrix shows dominance along the diagonal, which corresponds to in-distribution tests. From this diagonal, we infer that the temple vibration pickup and the throat microphone pose the greatest challenge for ASR in quiet conditions, given their limited bandwidth. A comparison between models trained on airborne and temple vibration sensors reveals a significant trend: the temple vibration pickup demonstrates enhanced robustness when tested on other sensors data. This trend can be attributed to the model’s reliance primarily on low-frequency information, which are present on all sensors. With regard to the in-ear microphones, there is a possibility that a single model can be developed in future studies, since they are highly similar and exhibit minimal performance loss when cross-tested. At last, the performance of the forehead accelerometer is approaching that of the in-ear microphone, despite the larger bandwidth. This may be attributed to the considerable variance in the coherence function and the lower signal-to-self-noise ratio.
Figure 6: PER of the finetuned audio phonemizer. Model trained and tested on the speech-clean subset

### v4

@@ III-B Textual data
The text utilized for participant readings originates from the French Wikipedia subset of Common Voice [62]. This textual data was further filtered to produce a simplified dataset with a minimum number of textual and phonemic symbols and to minimize pronunciation uncertainty. To create this simplified dataset, we applied textual filters that excluded most proper names, retaining only common first names, French town or region names, and country names. Moreover, utterances containing numbers, Greek letters, mathematical symbols, or syntactic errors were carefully removed, resulting in a more straightforward dataset that is suitable for accurate pronunciation analysis. Once the unwanted sentences were removed, we applied common text normalization operations, such as lowercasing text and removing punctuation except for apostrophes. From this normalized text, we also generated the corresponding phonetic transcription using the International Phonetic Alphabet (IPA), which contains 33 phonemes in French. This operation was done with the text phonemizer of the CoML lab [63]. For some words, the phonemizer produced some language switches that we wanted to avoid, to preserve the minimal French phonetic alphabet. A phase of manual phonetic transcription was then necessary for these unrecognized words, such as ”algorithme”.
@@ III-D Recording protocol
  • speechless-noisy: For a period of 2 minutes and 24 seconds, the subject is required to remain silent yet free to move, swallow, and breathe naturally in a noisy environment created from the AudioSet samples described in subsection III-C. Those samples are played in a spatialization sphere equipped with 56 loudspeakers surrounding the subject. The objective of this phase is to gather realistic background noises that will be combined with the speech-clean recordings to maintain a clean reference.
@@ III-E Post processing
The filtering process comprises two main stages. In the first stage, indicators for evaluating audio capture quality are extracted (Figure 2(a)). In the second stage, these indicators are used to derive a set of filters (Figure 2(b)). The whisper-timestamped model333available at https://github.com/linto-ai/whisper-timestamped constitutes a key component of our indicator extraction process. This model is a speech-to-text engine based on Whisper [29] that applies dynamic time warping [67] to cross-attention weights. By aligning the transcription with the raw waveform, accurate prediction of word timestamps in speech segments can be used to identify sequences containing vocalizations. The indicators we are using are listed below:
  • $\mathrm{PER}[j]$: Phoneme Error Rate is computed for each sample $j\in[1,J]$ using the Levenshtein distance between the phonetic labeled transcription and the phonetic transcription derived from Whisper’s output. Note that we used the exact same phonemization procedure as described in subsection III-C.
@@ III-F Signal processing analysis
Figure 3: Coherence functions of body-conducted sensors in quiet conditions
Each sensor reveals a distinct bandwidth size that the figure legend follows from left to right in decreasing order. The temple sensor exhibits the smallest bandwidth, barely recording any speech signal above 1500 Hz. The laryngophone follows, with coherence dropping below 0.25 at 2000 Hz. The soft and rigid in-ear sensors exhibit larger bandwidths yet display clear antiresonance frequencies. The soft in-ear bandwidth is slightly larger than its rigid in-ear counterpart, which may be due to a less effective acoustic seal. Finally, the forehead accelerometer offers the largest bandwidth but does not effectively filter out external noise pollution. As discussed in Section IV, this bandwidth is closely correlated with the performance of deep learning models on speech intelligibility, quality, speaker verification, and speech-to-phonemes.
@@ IV Using Vibravox
The Vibravox dataset offers considerable potential across various applications, particularly in improving communication in noisy environments. The speech-noisy subset is of particular interest for showcasing the noise-resilient capabilities of BCMs. However, this subset primarily serves testing purposes, as it lacks a clean reference under such conditions. As a consequence, the majority of the dataset is contained within the speech-clean subset, where the airborne microphone serves as the reference. In order to simulate real-world noisy scenarios while retaining a clean reference, data augmentation techniques can be applied by adding audio segments of the speechless-noisy subset with BCM utterances. Furthermore, the Vibravox dataset displays potential for applications such as developing voice commands and monitoring vital signs, including heart rate, particularly when utilizing the speechless-clean subset. In the following, we propose three classic tasks in automatic speech processing to investigate the distinctive characteristics of BCMs, establishing baselines for each task.
@@ IV-A Speech Enhancement
The high impedance of BCMs enables them to focus on the wearer’s voice by rejecting ambient noise, but this also affects their ability to capture mid and high frequency components due to the intrinsic low-pass characteristics of the biological pathway of vibrations between the vocal cords and the body-conducted audio sensors. The absence of these frequencies in captured speech directly impacts the perceived quality, intelligibility, and distinctiveness. This highlights the necessity for a post-capture enhancement algorithm.
@@ IV-A1 Model
We adopted the configurable EBEN [17] approach, enriched with recent trends in the audio field. This GAN [70] architecture offers several advantages. In particular, the lightweight design of the 1D U-net-like generator is compatible with integration in mobile and wearable systems, where memory is limited and real-time is a priority. Additionally, the PQMF bank [71] offers the advantage of manipulating frequency content without the need to consider phase or manage complex values, as it operates directly on waveform signals. It is important to note that the limitations of GANs, including the difficulty of reaching a Nash equilibrium, the lengthy training period, and the additional computational cost associated with the discriminators, do not affect the model once it has been deployed.
@@ IV-A2 Architecture configuration
Separate EBEN models were trained for each sensor in the Vibravox dataset, using sensor-specific hyperparameters detailed in Table IV. The original number of subbands was maintained at a value of $M=4$ for each sensor due to the significant variability in intra-sensor bandwidth among participants, rendering precise spectrogram slicing impractical. Furthermore, a smaller number of subbands allows for a more permissive design of the passband filters with a softer slope for the same resulting aliasing, thus shortening the PQMF kernel length, which is useful for reducing the algorithmic latency. The number of subbands provided to the model, denoted as $P$, was tailored to match the bandwidth of each sensor. For example, the temple vibration pickup captures signals within the 0-2 kHz range, while the forehead accelerometer exhibits good coherence with the reference microphone up to 8 kHz. Consequently, after resampling the signals to 16 kHz for training, only the first band was retained for the temple vibration pickup, while all bands were retained for the accelerometer. In addition, the last configurable parameter, $Q$, which represents the number of bands fed to the PQMF discriminator, has been modified to match $M$ at 4 for all models. It was observed that even when the audio sensors capture information within a specific frequency range – including the corresponding subbands in the discriminator input – it is beneficial for denoising minor artifacts present in these initial bands. In the case of speech captured in noisy conditions, even though body conduction sensors are robust to external acoustic pollution, some noise is still captured, providing an additional reason to focus discriminators work on every subband. This minor modification necessitated an adjustment to the number of channels in the PQMF discriminators in order to align with the divisibility constraints of the grouped convolution.
Figure 4: Spectrograms of signals recorded by the different Vibravox audio sensors in the speech-clean subset and their corresponding EBEN-enhanced version
@@ IV-A3 Training procedure
In comparison to the original implementation proposed in [17], we have incorporated a spectral loss $\mathcal{L}_{\mathcal{G}}^{spec}$ into the generator objective. This modification proved to be effective in stabilizing the training process. This loss incorporates the same FFT, hop size, and window length parameters as described in [72]. Implementation is taken from [73]. In contrast to some speech enhancement papers [74, 75], we did not include a reconstructive loss in the temporal domain due to the peculiarities of body-conduction speech. In fact, complex phase shifts occur : attempting to align the signals to the reference at the sample level degraded performance. We retained the original EBEN implementation’s use of the other two adversarial and a feature matching losses, respectively noted $\mathcal{L}_{\mathcal{G}}^{adv}$ and $\mathcal{L}_{\mathcal{G}}^{feat}$. Also note that a slight adjustment to $\mathcal{L}_{\mathcal{G}}^{feat}$ has been done by normalizing it using the enhanced feature norm. The training strategy remained consistent across speech-clean and speech-noisy experiments, with a constant learning rate of $3e^{-4}$. To ensure reproducibility of experiments, the code has been made available on the GitHub repository https://github.com/jhauret/vibravox, which makes use of the following libraries [76, 77, 78, 79].
@@ IV-A4 Experiments in quiet conditions
For this initial set of experiments, the goal is to evaluate the performance of bandwidth extension in quiet conditions by training EBEN models on the speech-clean subset. A total of 500 epochs were required for convergence on the Vibravox dataset. To address the relatively small amount of data, light data augmentation was applied, including speed perturbation, pitch shifting, and time masking, all implemented in the Torchaudio library [80]. The correlation analysis presented in the configurable EBEN article [17] revealed that STOI [81, 82] and Noresqa-MOS (N-MOS) [83] were the metrics with the highest correlation with MUSHRA studies in the context of speech enhancement on BCMs simulated data. Consequently, these two metrics were retained for inclusion in Table IV, which presents the results of the bandwidth extension for each sensor in quiet conditions. The results reveal a clear relationship between the metrics value and the sensors’ available bandwidth. Yet, both STOI and N-MOS performance appear to hit a glass ceiling, plateauing at 0.88 and 4.3, respectively. This saturation could stem from the challenge of extracting fine details from body-conducted signals. While striving to produce authentic samples, the generator only picks one plausible clean signal out of all the possible ones, typical of one-to-many problems [84]. The spectrograms in Figure 4 provide a visual representation of the processing.
@@
Sensor Configuration STOI Noresqa-MOS
@@ IV-A5 Experiments in noisy conditions
After analyzing speech enhancement in quiet conditions, the next focus is a joint denoising and bandwidth extension task for recordings in noisy environments. To this end, the model is tested on the speech-noisy subset; however, this subset cannot be used for training due to the absence of a clean reference. For training, the speechless-noisy subset is leveraged and dynamically mixed with the speech-clean subset. Several configurations were tested, including retraining the EBEN models from scratch for 500 epochs or initializing from the previously trained speech-clean models, which converged after no more than 200 epochs for microphones with the largest domain gap. Additionally, speech-clean models were directly tested on speech-noisy. Hyperparameters of the training strategy were kept consistent with quiet condition experiments, with the same values for M, P, and Q across all microphones.
Prior to testing on speech-noisy, the model performance was first evaluated on the validation and test splits of the synthetically mixed speech-clean+speechless-noisy data. When testing on speech-noisy, no significant differences were observed in the Noresqa-MOS scores, reflecting a realistic modeling of the training data. When using speech-noisy for evaluating the models, the STOI – which is a reference-based metric – was replaced with the reference-free TorchSquim-STOI [85], which provides an estimate of the same quantity. Although direct comparison between STOI and TorchSquim-STOI is not possible, both metrics showed improvement during training, and the TorchSquim-STOI results remained consistent across both synthetic mixed and real noisy data. The final test results on speech-noisy are presented in Table V.
@@
Sensor Initialization Squim-STOI Noresqa-MOS
@@
The results reveal that microphones with the lowest signal-to-external-noise ratios benefited the most from the retrained strategies. Some microphones, such as the temple vibration pickup, were so insensitive to noise that the non-retrained speech-clean models outperformed them. The Noresqa-MOS scores for the raw signals also increased compared to the speech-clean subset. However, these Noresqa-MOS values cannot be directly compared, as in the quiet conditions, the metric was computed using matched signals, whereas in the noisy conditions, a random headset signal from the speech-clean was used for the calculation.
@@
As with the experiments conducted in quiet conditions, several spectrograms are presented in Figure 5. Synthetically mixed data were selected to visually assess the effect of the models, as both the noise and the clean target signals remain available. This figure highlights several noteworthy observations. Notably, the excessive noise detected in the soft in-ear microphone suggests an insertion defect for this particular sample of speechless-noisy, whereas the rigid microphone appears to provide better acoustic sealing. However, such defects occur just as often in both microphone types and are absent in most cases. Additionally, EBEN models trained for bandwidth extension in quiet conditions exhibit limited noise suppression. In contrast, as shown in line (c), noise-robust EBEN models effectively suppress noise, with the difference particularly pronounced for forehead and in-ear microphones. For throat and temple microphones, which are inherently more resistant to noise, the domain gap between quiet and noisy conditions remains small, enabling the speech-clean EBEN models to maintain strong performance. Achieving comparable denoising performance from the noisy headset using conventional denoising networks would be highly challenging.
@@ IV-B2 Experiments in quiet conditions
For this first set of experiments, we finetuned the wav2vec2.0 models with a constant learning rate of $1e^{-5}$ on the speech-clean subset of the Vibravox dataset for 10 epochs, minimizing the CTC loss [88]. We then tested each fine-tuned model not only on its corresponding test set but also on every other audio sensor, with the PER (Phoneme Error Rate) results shown in Figure 6. As expected, the matrix shows dominance along the diagonal, which corresponds to in-distribution tests. From this diagonal, we infer that the temple vibration pickup and the throat microphone pose the greatest challenge for ASR in quiet conditions, given their limited bandwidth. A comparison between models trained on airborne and temple vibration sensors reveals a significant trend: the temple vibration pickup demonstrates enhanced robustness when tested on other sensor data. This trend can be attributed to the model’s reliance primarily on low-frequency information, which is present on all sensors. Regarding the in-ear microphones, there is a possibility that a single model can be developed in future studies, since they are highly similar and exhibit minimal performance loss when cross-tested. At last, the performance of the forehead accelerometer is approaching that of the in-ear microphone, despite the larger bandwidth. This may be attributed to the considerable variance in the coherence function and the lower signal-to-self-noise ratio.
Figure 6: PER of the finetuned audio phonemizers. Models are trained and tested on the speech-clean subset